OpenStudy (fanduekisses):

Find the sum of the first 9 terms of the sequence. 2, -7, -16, -25, ...

OpenStudy (fanduekisses):

do I have to find the last term first?

OpenStudy (fanduekisses):

l=a + (n-1)d , where l= last term, a= first term, n= # of terms, d= difference

OpenStudy (fanduekisses):

@jdoe0001

OpenStudy (anonymous):

dont you just find the pattern, find 1st 9 num, then add?

OpenStudy (fanduekisses):

help, no the

OpenStudy (fanduekisses):

sigma

OpenStudy (fanduekisses):

\[\Sigma \]

OpenStudy (anonymous):

This is obviously an arithmetic series. To find the common difference, subtract the first term from the second term... -7 - 2 = - 9....so the common difference is -9. Now before we find the sum, we need to find the 9th term an = a1 + (n-1) * d n = term we want to find = 9 a1 = first term = 2 d = common difference = -9 now we sub a9 = 2 + (9-1) * -9 a9 = 2 + 8 * -9 a9 = 2 - 72 a9 = - 70 now that we know the 9th term we will then find the sum sn = n(a1 + an) / 2 n = terms we need to add = 9 a1 = first term = 2 a9 = -70 now we sub s9 = 9 (2 - 70) / 2 s9 = 9 (-68) / 2 s9 = 9 * - 34 s9 = - 306 so the sum of the first 9 terms is -306

OpenStudy (fanduekisses):

yay thanks<3